Shift summary for owl 8/23/99

Jason Hamm (jhamm@physics.arizona.edu)
Mon, 23 Aug 1999 05:00:56 -0700 (MST)


Shift Summary for owl shift, 08/23/99

On Shift: S. Ledovsky, J. Hamm

Hours of beam available at all intensities: 7
Hours of beam available at nominal intensity: 7
Hours of E832 data (special runs not included): 6.75
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time: Changing tapes
========================================================================
Runs ended this shift:

14301 - 34.5 K vac 2pi0

Tapes brought to FCC this shift: QK(S,C,K)309-314, QKB108-109
QKA016, QKNL25, QKN027, QKLA14
========================================================================
Accesses during shift: none
========================================================================
Overview of shift: Unbelievably smooth running at ~6.8 E12
========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
3:50 Ktev3c 91% full. We are directed to delete the
onsplit caches corresponding to the tapes listed
above.

4:15 Onsplit requests QKA017 on ktev3

4:25 Tapes fill up, run 14301 ends. We fresh start, as
requested by the previous shift.

4:40 Start Beam_832 run 14302

========================================================================
Follow up Notes on Previous Shift
========================================================================
Items to Follow up on Next Shift

Not a single ktevana error all night. Keep the streak going!
CsI 166 had comparator problems in 14301 (Freq=6).